# The ML Supply Chain in the Era of Software 2.0: Lessons Learned from Hugging Face

### v1

@@ 5.3.1. Lengths of model supply chains
We define a lineage chain as a path from some root base model (i.e., one with no model dependencies) to a final sink node model (i.e., one without any model dependents). We examine 53,151 lineage chains for models that declare at least one base model, including cases where multiple chains lead to the same model. Of these, the average chain length to reach that model is 6.2 models. The most frequent chain length is of three models (12,480 chains), and, as shown in Figure 6, the first quartile is 3 models, the median is 4 models, and the third quartile is 9 models. The longest chain, beginning with cohereforai/c4ai-command-r-v01 and ending with Citaman/command-r-1-layer, contains 40 models. Excluding the first model, all models in this chain are owned by Citaman and incrementally count down from 39. That is, the models in the chain range from Citaman/command-r-39-layer down to Citaman/command-r-1-layer, suggesting that each model may represent an incremental improvement or an updated version.
Figure 5. Model Chain Lengths
Figure 6. Number of Licenses in Model Chains
@@ 5.3.2. Licenses present in model supply chains
There are typically relatively few parent/child license differences between models in these chains, with an average of 1.9 distinct licenses in a given chain, including unknown or undeclared licenses. As Figure 6 illustrates, the first quartile is just 1 license, and both the median and third quartile are 2 licenses. A chain most commonly contains 2 licenses (24,069 chains), and the highest number of licenses observed was 6, occurring in 17 different chains. A relatively small number of parent/child license differences simplifies the task of managing license compliance. However, these differences within a lineage necessitate examination of license compatibility, as well as a possible need to find mitigations for items in model chains when required by the license of a base model.
Figure 7. Number of Times the Owner of a Chain’s Final Model Appears in that Chain
@@ 5.3.3. Model ownership
We also observe that developers using Hugging Face often build off each other’s work rather than off their own models. The owners of nodes at the end of a chain appeared within that node’s chain only 1.4 times on average, showing that base models are frequently sourced from the community and supporting the idea that a common approach to training a model involves building off of previous work. In another interpretation, it could be that model trainers do not fill in base model information when building on their own work at least some of the time. Future work involving discussions with developers could shed light on the reasons behind this phenomenon.
In 82.9% of chains (43,507), the owner of the final model in the chain owned no other model in that chain. In the maximum case, the chain contained 39 models from the final model’s owner. Illustrating this further, as shown in Figure 7, the first quartile, median, and third quartile are all a single occurrence of the final model’s owner in a given chain: only the final model itself.
Based on our observations, a given owner publishes relatively few models to Hugging Face, owning, on average, 4 models in our dataset, though the data is skewed heavily toward relatively few highly-prolific owners. Of 190,136 distinct model owners in our dataset, 61.5% owned just one public model (117,016), while the most prolific account owned 4,610 models. The first quartile and median are both just 1 owned model, and the third quartile is 2 models. Notably, the large players mentioned above, including Meta and OpenAI, are not present in the top-10 most prolific model owners: those spots instead go to smaller stakeholders in the AI market. The top-ten model owners on Hugging Face are shown in Table 11 and the top-ten dataset owners in Table 11. There is only one overlap between the two lists.
Table 10. Top-10 model owners on Hugging Face
Owner Model Count
mradermacher 4610
thebloke 3863
huggingtweets 3830
tkcho 3810
jeevesh8 3723
prunaai 3411
lonestriker 3241
sail-rvc 3182
larryaidraw 3046
cyberharem 2978
Table 11. Top-10 dataset owners on Hugging Face
Owner Dataset Count
open-llm-leaderboard-old 7059
cyberharem 3384
autoevaluate 3232
recag 2345
liuyanchen1015 1558
adapterocean 810
tyzhu 807
joey234 705
lots-of-loras 500
results-sd-v1-5-sd-v2-1-if-v1-0-karlo 462
@@ 6. RQ2: licensing of models and datasets on Hugging Face
The majority of models (62.2%) and datasets (72.4%) did not declare licensing information in a machine-readable way (see Section 3.3). To better illustrate how models and datasets are licensed, we organize the licenses that we observed into six categories. The categories are intended to be indicative of the origin or purpose of the licenses (whether the licenses were designed for open source software, for ML components specifically, or other purposes) and are not intended to convey any legal characteristics, particularly given the uncertainty involved in applying these licenses in the ML context. We define the following categories:
@@ 6.1. Most common licenses
Table 12. Top-10 licenses for models.
@@
Table 13. Top-10 licenses for datasets.
@@
cc-by-sa-4.0 CC 1,256 2.6%
cc-by-nc-4.0 CC 1,014 2.1%
cc-by-nc-sa-4.0 CC 946 2.0%
cc0-1.0 CC 799 1.7%
There were 71 unique declared licenses for models and 70 for datasets. Tables 13 and 13 provide an overview of the most frequently observed licenses.
We classify licenses by the classes defined above. A breakdown of the prevalence of these classes can be found in Table 14. Both models (62.5%) and datasets (58%) are most likely to be licensed under approved OSS licenses. This is particularly interesting for datasets, where we would suspect to see mostly CC or Data licenses, as datasets are not “software” in a strict sense of the word. This may reflect that Hugging Face developers view models and datasets more similarly to programs than to data or that developers rely on OSS licenses because they are already familiar with their terms. However, the relationships between these preexisting licenses and these new types of components in the ML supply chain, as well as the implications of licensing such components with such licenses, are currently unknown. Further research is needed to understand the reasoning behind these licensing choices, as well as to understand the interactions between many of these licenses and ML components.
@@
Total Licensed 287,143 100.0% Total Licensed 48356 100.0%
@@
Both models and datasets are distributed with (i) OSS licenses specific to software (both restrictive and permissive), (ii) Creative Commons (CC) licenses, that are not software-specific yet are being used for software (e.g., Stack Overflow’s adoption of CC licenses), and ML-specific licenses. For models, our findings are consistent with those of Pepe et al. (Pepe et al., 2024), showing that permissive OSS licenses such as Apache-2.0 and MIT are popular in this space. This is not surprising, as such licenses are also among the most popular for open-source projects (Vidal, 2023; Vendome, 2015) and allow commercial/closed-source exploitation. We note that, in general, the top licenses adopted on Hugging Face are traditionally understood to be more permissive, even if their exact application in this novel context is not fully understood.
@@
We also note the adoption (for models, but especially for datasets) of different variants of the CC license, including a permissive one (CC-BY), restrictive ones (CC-BY-SA and CC-BY-NC-SA), and licenses limiting non-commercial use only (CC-BY-NC).
@@
Besides OSS licenses, we see a large proportion of different kinds of ML-specific licenses. These include both licenses originating from open-source initiatives (openrail, creativeml-openrail-m), and licenses originating from companies (e.g., Meta’s llama2 and llama3 licenses). Such licenses tend to differ from typical redistribution terms established by software licenses by, for example, introducing “behavioral” constraints; requiring responsible model usage; and prohibiting adoption for harmful or unethical uses or uses that do not take models’ limitations into account. Likewise, the llama2 and llama3 licenses impose limitations related to commercial use, use in products or services having more than 700 million monthly active users, and use to train other models, unless such models are redistributed, under the same licenses, as derivative works of llama models.
@@
Last but not least, we note the increased prevalence of Hugging Face’s “Other” license which was ranked 7th in Pepe et al. (Pepe et al., 2024) but has moved to the 5th spot since. In fact, 3.8% of datasets and 5.7% of models are licensed with some “Other” license. Prior work has found that such license proliferation can make license compliance tasks more time-consuming and difficult, since compliance teams are no longer dealing with known quantities (Wintersgill et al., 2024).

### v2

@@ 6.4.1. Lengths of model supply chains
We define a lineage chain as a path from some root base model (i.e., one with no model dependencies) to a final sink node model (i.e., one without any model dependents). We found and examine 53,151 lineage chains for models that declare at least one base model, including cases where multiple chains lead to the same model. Of these, the average chain length to reach that model is 6.2 models. The most frequent chain length is three models (12,480 chains); the minimum is 2 models, the first quartile is 3, the median is 4, the third quartile is 9, and the maximum is 40.
@@ 6.4.2. Licenses present in model supply chains
There are typically relatively few parent/child license differences between models in these chains. We remind the reader than a parent/child license difference does not necessitate an incompatibility, but instead demonstrates that some licensing decisions (or lack thereof in the case of neglecting to license a derivative work) were made along the chain. There was an average of 1.9 distinct licenses in a given chain, including unknown or undeclared licenses. The minimum and first quartile are both just 1 license, both the median and third quartile are 2 licenses, and the maximum is 6 licenses. A chain most commonly contains 2 licenses (24,069 chains), and the highest number of licenses observed was 6, occurring in 17 different chains. A relatively small number of parent/child license differences simplifies the task of managing license compliance. However, these differences within a lineage necessitate examination of license compatibility, as well as a possible need to find mitigations for items in model chains when required by the license of a base model.
@@ 6.4.3. Model ownership
We also observe that when choosing models to build off of, developers using Hugging Face often use models from the community rather than using their own previous work as a base. The owners of nodes at the end of a chain appeared within that node’s chain only 1.4 times on average, showing that base models are frequently sourced from the community and supporting the idea that a common approach to training a model involves building on previous work. Alternatively, it could be that model trainers do not provide base model information when building on their own work at least some of the time. Future work involving discussions with developers could shed light on the reasons behind this phenomenon.
In 82.9% of chains (43,507), the owner of the final model in the chain owned no other model in that chain. In the maximum case, the chain contained 39 models from the final model’s owner. The minimum, first quartile, median, and third quartile are all a single occurrence of the final model’s owner in a given chain: only the final model itself.
Based on our observations, a given owner publishes relatively few models to Hugging Face, owning, on average, 4 models in our dataset, though the data is skewed heavily toward relatively few highly-prolific owners. Of 190,136 distinct model owners in our dataset, 61.5% owned just one public model (117,016), while the most prolific account owned 4,610 models. The first quartile and median are both just 1 owned model, and the third quartile is 2 models. Notably, the large players mentioned above, including Meta and OpenAI, are not present in the top 10 most prolific model owners: those spots instead go to smaller stakeholders in the AI market. The top ten model and dataset owners on Hugging Face are shown in Table 12. There is only one overlap between the two lists.
Table 12. Top 10 model and dataset owners on Hugging Face
Models Datasets
Owner Model Count Owner Dataset Count
mradermacher 4,610 open-llm-leaderboard-old 7,059
thebloke 3,863 cyberharem 3,384
huggingtweets 3,830 autoevaluate 3,232
tkcho 3,810 recag 2,345
jeevesh8 3,723 liuyanchen1015 1,558
prunaai 3,411 adapterocean 810
lonestriker 3,241 tyzhu 807
sail-rvc 3,182 joey234 705
larryaidraw 3,046 lots-of-loras 500
cyberharem 2,978 results-sd-v1-5-sd-v2-1-if-v1-0-karlo 462
6.5. Summary
Our findings show that Hugging Face components rarely declare datasets, generally have short chains of related components, and when dependency relationships between components exist, licensing information usually stays the same from parent to child. This perhaps simplifies the task of managing dependencies and licensing, but largely by virtue of the fact that relevant supply chain information appears to be missing in many such cases.
@@ 7. RQ2: licensing of models and datasets on Hugging Face
Only 37.8% of models and 27.6% of datasets declared licensing information in a machine-readable way (see Section 4.3). In this section, we investigate this licensing information to better understand how datasets and models are licensed in the ML supply chain. We begin by organizing the licenses that we observed into six categories, intended to be indicative of the origin or purpose of the licenses (whether the licenses were designed for open source software, for ML components specifically, or for other purposes) and are not intended to convey any legal characteristics, particularly given the uncertainty involved in applying these licenses in the ML context. These categories are not aimed at resolving compliance issues but instead demonstrate the complex and diverse landscape of licenses in the ML supply chain. We define the following categories:
@@ 7.1. Most common licenses
Table 13. Top 10 licenses for models.
@@
Table 14. Top 10 licenses for datasets.
@@
$c$-by-sa-4.0 $C$ 1,256 2.6%
$c$-by-nc-4.0 $C$ 1,014 2.1%
$c$-by-nc-sa-4.0 $C$ 946 2.0%
cc0-1.0 $C$ 799 1.7%
There were 71 unique declared licenses for models and 70 for datasets. Tables 14 and 14 provide an overview of the ten most frequently observed licenses.
We classify licenses by the categories defined above. A breakdown of the prevalence of these classes can be found in Table 15. Both models (62.5%) and datasets (58%) were most often licensed under approved OSS licenses. This is particularly interesting with respect to datasets, as we would expect to see primarily CC or Data licenses, as datasets are not “software” in a strict sense of the word. This may reflect that Hugging Face developers view models and datasets more similarly to programs than to data or that developers rely on OSS licenses because they are already familiar with their terms. However, as prior work indicates (Duan et al., 2024), the relationships between these preexisting licenses and these new types of components in the ML supply chain, as well as the legal and other implications of licensing such components with such licenses, are currently unknown. Further research is needed to understand the reasoning behind these licensing choices, as well as to understand the interactions between many of these licenses and ML components.
@@
Total Licensed 287,143 100.0% Total Licensed 48,356 100.0%
@@
Both models and datasets are distributed with (i) OSS licenses specific to software (both restrictive and permissive), (ii) Creative Commons (CC) licenses, which are not software-specific yet are being used for software (e.g., Stack Overflow’s adoption of CC licenses), and ML-specific licenses. For models, our findings are consistent with those of Pepe et al. (Pepe et al., 2024), showing that permissive OSS licenses such as Apache-2.0 and MIT are popular in this space. This is not surprising, as such licenses are also among the most popular for open-source projects (Vidal, 2023; Vendome, 2015) and allow commercial/closed-source exploitation. We note that, in general, the top licenses adopted on Hugging Face are traditionally understood to be more permissive, even if their exact application in this novel context is not fully understood.
@@
We also note the adoption (for models, but especially for datasets) of different variants of the CC license, including a permissive license (CC-BY), restrictive licenses (CC-BY-SA and CC-BY-NC-SA), and licenses limiting non-commercial use only (CC-BY-NC).
@@
Besides OSS licenses, we see a large proportion of different kinds of ML-specific licenses. These include both licenses originating from open-source initiatives (e.g., openrail and creativeml-openrail-m), and licenses originating from companies (e.g., Meta’s llama2 and llama3 licenses). Such licenses tend to have terms that differ from typical redistribution terms established by software licenses such as, for example, terms that require responsible model usage or prohibit adoption for harmful or unethical uses or uses that do not take models’ limitations into account. Likewise, the llama2 and llama3 licenses impose limitations related to commercial use, use in products or services having more than 700 million monthly active users, and use to train other models, unless such models are redistributed, under the same licenses, as derivative works of llama models.
@@
Lastly, we note the increased prevalence of Hugging Face’s “Other” license, which was ranked 7th in a study conducted by Pepe et al. (Pepe et al., 2024), but has since moved to the 5th spot since. In fact, 3.8% of datasets and 5.7% of models are licensed with some “Other” license.
7.2. Exploration of the “Other” License
In our entire dataset, 5.7% of licensed models were labeled as being under some “Other” license. This makes the “Other” category the fifth most common licensing choice for models on Hugging Face, representing over sixteen thousand models. Non-standard, custom/vanity licenses can complicate compliance processes for organizations and individuals, given that the documents no longer represent a known quantity (Wintersgill et al., 2024). To assess the nature and potential impact of these licenses, we investigate these “Other” licenses in further detail.
7.2.1. Manual analysis of models under the “Other” license
We examine the licenses of the top 25 “Other”-licended models, sorted by downloads, as well as those of 25 additional, randomly sampled “Other”-licensed models from the total population. Our assumption is that the popularity of models in the high-downloads group indicates a tendency toward high-quality documentation practices, allowing us to investigate a detailed and meaningful sample of projects licensed in this way while maintaining a reasonable sample size for manual analysis. We also select a random sample in an attempt to determine if the same licensing trends exist across the ecosystem, even for less popular models. One author manually and systematically reviewed the model pages of these 50 models to identify more specific licensing information. Cumulatively, these 50 models represent 13,945,199 downloads (0.8% of all model downloads). The results can be seen in Tables 16 and 17. (One of the models randomly sampled was no longer available on Hugging Face for analysis.)
Table 16. “Other” license information for Top 25 “Other”-licensed models
Model Name Resolved License Find Location
facebook/opt-125m OPT-175B LICENSE AGREEMENT LICENSE.md
stabilityai/sdxl-turbo sai-nc-community LICENSE.md
coqui/XTTS-v2 coqui-public-model-license LICENSE.txt
facebook/mask2former-swin-base-coco-panoptic None Available* None Available
briaai/RMBG-1.4 BRIA 2.0 on HuggingFace Model License Agreement
External Link
(connected to tag)
facebook/mask2former-swin-large-cityscapes-semantic None Available* None Available
apple/DFN5B-CLIP-ViT-H-14-378 apple-sample-code-license LICENSE
facebook/opt-1.3b OPT-175B LICENSE AGREEMENT LICENSE.md
facebook/mask2former-swin-large-ade-semantic None Available* None Available
Lykon/DreamShaper
CreativeML Open RAIL-M
Addendum*
External Link
(in model card)
Salesforce/instructblip-vicuna-7b None Available* None Available
SanctumAI/Meta-Llama-3-8B-Instruct-GGUF llama3 LICENSE
cagliostrolab/animagine-xl-3.1 faipl-1.0-sd
External Link
(connected to tag)
stabilityai/stable-video-diffusion-img2vid-xt stable-video-diffusion-community LICENSE.md
apple/mobilevit-small apple-sample-code-license In model card
jartine/gemma-2-27b-it-llamafile Gemma License In model card*
facebook/opt-350m OPT-175B LICENSE AGREEMENT LICENSE.md
Qwen/Qwen2-72B-Instruct-AWQ tongyi-qianwen LICENSE
Mozilla/Meta-Llama-3-70B-Instruct-llamafile Llama3 In tags*
digiplay/AbsoluteReality_v1.8.1 Absolute Reality License*
External Link
(in model card)
qwp4w3hyb/Qwen2-72B-Instruct-iMat-GGUF tongyi-qianwen In tags*
google/mobilenet_v1_0.75_192 None Available None Available
Qwen/Qwen2-72B-Instruct tongyi-qianwen LICENSE
mradermacher/QuartetAnemoi-70B-t0.0001-i1-GGUF None Available None Available
SanctumAI/Codestral-22B-v0.1-GGUF mnpl In tags*
* Information was ultimately inconsistent, unavailable, or changed since initial mining.
Table 17. “Other” license information for random sample of “Other”-licensed models
Model Name License Location
Shalie/EliraPendoraPonyXL faipl-1.0-sd External Link (connected to tag)
sukhoi37/sp_weshi None Available* None Available
Bobross69/Usagiyam None Available* None Available
baratk/barat Empty file LICENSE file
styalai/competition-math-phinetune-v1 None Available* None Available
divyanshu074/Text_Beat None Available* None Available
LoneStriker/WhiteRabbitNeo-33B-v1-3.0bpw-h6-exl2 deepseek External Link (connected to tag)
jiuhai/llama3-ift REMOVED* REMOVED*
KnutJaegersberg/Deacon-34B-qlora yi-license LICENSE file
second-state/Qwen1.5-1.8B-Chat-GGUF tongyi-qianwen-research LICENSE file
psychicfire/sample None Available* None Available
Workaholic/fake-pvc-style None Available* None Available
azi111/dolphin-2_2-yi-34b-465bpw-h8-exl2-cnen Empty file LICENSE file
dranger003/deepseek-coder-33b-instruct-iMat.GGUF deepseek External Link (connected to tag)
optimai/porne-0 None Available* None Available
cortecs/Meta-Llama-3-70B-Instruct-GPTQ-8b None Available* None Available
TheBloke/MergeMonster-13B-20231124-GPTQ Other / llama2 Model card
yzhuang/Meta-Llama-3-8B-Instruct_fictional_Chinese_v3 None Available* None Available
fasterinnerlooper/stable-code-3b None Available* None Available
DevQuasar/Qwen2-72B-Instruct-GGUF tongyi-qianwen External Link (connected to tag)
titan087/Llama3-70B-ShiningValiant2-exl2-4b llama3 External Link (connected to tag)
nulltella/phi-1_5-alcapa-51k-instruct None Available* None Available
Holarissun/mar13_gemma2b_aisft_gsm8k_rand_alpha0.9995-subset7000 None Available* None Available
raincandy-u/Quark-464M-v0.1.alpha tongyi-qianwen-research External Link (connected to tag)
TheBloke/Tess-34B-v1.4-GGUF yi-34b External Link (connected to tag)
* Information was ultimately inconsistent, unavailable, or changed since initial mining.
7.2.2. Top 25 “Other” licensed models
For the top 25 models, additional licensing information was most often found in a LICENSE file that was associated with the repository, but four (16%) linked to an external page that included more licensing information. There were four cases (16%), including three Facebook models, where a specific license could not be determined, but one could be inferred from a related GitHub repository, which often was not directly linked to the model card. We also observed a case, Lykon/DreamShaper, where the license that was specified in the model card did not match the external reference also provided.
Determining the licenses for some models required looking beyond the information provided on Hugging Face itself. For example, the README for the model digiplay/AbsoluteReality_v1.8.1 included a link to a model page on an external site (Civitai.com). On this page, the license is reported as CreativeML Open RAIL-M Addendum. Originally, clicking the link associated with the license would take the user to a page for the Absolute Reality License (a confusing license mismatch), but in the time since, the link has been corrected to point to the full text of the CreativeML Open RAIL-M Addendum.
We can also see there were three instances where a model reported an “Other” license, but the actual license specified was in Hugging Face’s recognized license list (either llama3 or gemma). This could be the result of the model owner uploading their model before the license was officially recognized or a general unfamiliarity on the uploader’s part with the available license options. In either case, this particular error further underscores the manual and error-prone documentation processes on Hugging Face.
Somewhat surprisingly, even after systematic manual documentation searches (including web searches for associated GitHub repositories), there were two models in the top 25 “Other” licensed models (8%) where the specific license could not be identified. From a practical standpoint, these models might as well not be licensed at all, since users cannot comply with unavailable terms and conditions.
Interestingly, two of the models that had no licensing information included in their documentation, Salesforce/instructblip-vicuna-7b and google/mobilenet_v1_0.75_192, actually had model cards produced by the Hugging Face team.
7.2.3. Random sample of 25 Other-licensed models
Next, we examined an additional random sample of 25 models that were reportedly under an “Other”’ license. This would allow us to determine if the licensing practices followed by the top 25 are representative of the ecosystem. As one might expect, we found that identifying the licenses for many of these models was difficult. Ten models (40%) had model cards that were void of all text. One additional model had some information in its model card but none that pertained to licensing. Two more models had LICENSE files in their repositories, but the files were empty. Another model, Meta-Llama-3-70B-Instruct-GPTQ-8b, provided no further licensing information, but given its name, one might infer it was under the Llama3 license.
Some of the models, however, did include more specific licensing information. Six (24%) had links to external references included in their metadata. One model, while labeled as licensed under “Other”, was actually under the Llama3 license. Two models had LICENSE files that provided additional license information.
In the text of the model card for TheBloke/MergeMonster-13B-20231124-GPTQ, we observed the model owner grappling with certain ambiguities and uncertainties surrounding the “Other” license: “The creator of the source model has listed its license as other, and this quantization has therefore used that same license. As this model is based on Llama 2, it is also subject to the Meta Llama 2 license terms, and the license files for that are additionally included. It should therefore be considered as being claimed to be licensed under both licenses. I contacted Hugging Face for clarification on dual licensing but they do not yet have an official position. Should this change, or should Meta provide any feedback on this situation, I will update this section accordingly. In the meantime, any questions regarding licensing, and in particular how these two licenses might interact, should be directed to the original model repository.” Here, we see that properly maintaining license compliance can be a difficult and confusing task, particularly when all relevant information is not made readily available. It is also important to note that since the original model being quantized was itself a derivative of a Llama2 model, this model uploader has restored the Llama2 license to remain compliant.
7.2.4. Automated analysis of models under the Other license
Having determined from the manual review of 50 models reportedly under the “Other” license how and where to find more specific licensing information, we next moved to an automated approach that could provide insights about the most common “Other” licenses in the ecosystem. We note that since this analysis required access to more data than was originally pulled for the rest of the dataset described in this paper, we conducted a separate smaller mining on February 1, 2025. Specifically, we pulled the top ten thousand models (sorted by downloads) that reported they were under the “Other” license. We acknowledge that the results of this subsequent mining may not perfectly align with the dataset presented in the rest of this work, but nonetheless, we believe it is worthwhile to explore the phenomenon of the “Other” license in more depth.
Table 18. “Other” licenses reported in CardData attribute in Top 10,000 “Other”-licensed models
License Name Count Percentage
faipl-1.0-sd 2346 23%
flux-1-dev-non-commercial-license 1172 12%
llama3 269 3%
deepseek 206 2%
tongyi-qianwen 201 2%
Table 18 shows the top 5 specific licenses that were reported in the CardData attribute. Over a third of models (36%) did not provide additional licensing information in their CardData. Three percent of models are actually under the llama3 license, which is in the recognized license list on Hugging Face (Face, 2024a). As mentioned before, this suggests that either these models were uploaded before llama3 was added to the official list and have not been updated or that the model owners are unfamiliar with the licenses available in the official list. In either case, feasible tooling solutions can be readily imagined to resolve these problems.
There is a further lack of standardization in the licenses reported as “Other”. For example, 206 models report deepseek and an additional 85 report the deepseek-license, though these are presumably the same license. As previously mentioned, 269 models erroneously report llama3 as their “Other” license, but an additional 27 report being under the llama-3 license. We see 110 models under the gemma-terms-of-use, even though the proper identifier for this model is just gemma (Face, 2024a). All of this further underscores the problems inherent in providing developers with a free-text field instead of assisted tool support that suggests known licenses.